OpenStudy (anonymous):

|x-7\3|=6 absoulute value equations

OpenStudy (anonymous):

\[|\frac{x-7}{3}|=6\] or \[|x-\frac{7}{3}|=6\]?

OpenStudy (anonymous):

the second one @satellite73

OpenStudy (anonymous):

help @shamil98

OpenStudy (anonymous):

solve two equations \[x-\frac{7}{3}=6\]and \[x-\frac{7}{3}=-6\] both take one step, adding \(\frac{7}{3}\) to both sides

OpenStudy (shamil98):

Convert 7/3 to a decimal 3.5 then add it to 6 and -6 x = -6 + 3.5 x = -2.5 x = 6 + 3.5 x= 9